A couple invested part of P2,100,000 at 4% and the remainder at 7%. Their annual income from these two investments was equivalent to a return at 6% on the entire sum invested. how much money was invested at each rate?

Answer by mananth(16946)

Part I 4.00% per annum ------------- Amount invested =x
Part II 7.00% per annum ------------ Amount invested = y
2100000
6% of 21,000,000 = 126,000
Interest----- 126000.00

Part I 4.00% per annum ---x
Part II 7.00% per annum ---y
Total investment
x + 1 y= 2100000 -------------1
Interest on both investments
4.00% x + 7.00% y= 126000
Multiply by 100
4 x + 7 y= 12600000.00 --------2
Multiply (1) by -4
we get
-4 x -4 y= -8400000.00
Add this to (2)
0 x 3 y= 4200000
divide by 3
y = 1400000
Part I 4.00% $ 700000
Part II 7.00% $ 1400000

CHECK
700000 --------- 4.00% ------- 28000.00
1400000 ------------- 7.00% ------- 98000.00
Total -------------------- 126000.00

m.ananth@hotmail.ca